# Rethinking Invariance Regularization in Adversarial Training to Improve Robustness-Accuracy Trade-off

### v1

@@ Abstract
Although adversarial training has been the state-of-the-art approach to defend against adversarial examples (AEs), they suffer from a robustness-accuracy trade-off. In this work, we revisit representation-based invariance regularization to learn discriminative yet adversarially invariant representations, aiming to mitigate this trade-off. We empirically identify two key issues hindering invariance regularization: (1) a “gradient conflict” between invariance loss and classification objectives, indicating the existence of “collapsing solutions,” and (2) the mixture distribution problem arising from diverged distributions of clean and adversarial inputs. To address these issues, we propose Asymmetrically Representation-regularized Adversarial Training (AR-AT), which incorporates a stop-gradient operation and a predictor in the invariance loss to avoid “collapsing solutions,” inspired by a recent non-contrastive self-supervised learning approach, and a split-BatchNorm (BN) structure to resolve the mixture distribution problem. Our method significantly improves the robustness-accuracy trade-off by learning adversarially invariant representations without sacrificing discriminative power. Furthermore, we discuss the relevance of our findings to knowledge-distillation-based defense methods, contributing to a deeper understanding of their relative successes.
Deep Learning, Computer Vision, Adversarial Robustness, Trustworthiness, ICML
@@ 1 Introduction
(a) Naive invariance regularization
(b) Our approach: Asymmetrically Representation-regularized Adversarial Training (AR-AT)
Figure 1: Illustration of our approach. (a) Naive invariance regularization suffers from a “gradient conflict” between invariance loss and classification objectives, and a large distribution gap between clean and adversarial inputs. (b) Our approach uses an asymmetric structure with a stop-gradient operation and a predictor to avoid “gradient conflict”, and a split-BatchNorm (BN) to maintain consistent batch statistics during training.
@@
Computer vision models based on deep neural networks (DNNs) achieve remarkable performance in various tasks. However, they are vulnerable to adversarial examples (AEs) (Szegedy et al., 2014; Goodfellow et al., 2015), which are carefully perturbed inputs to fool DNNs. Since AEs can fool DNNs without affecting human perception, they are a severe threat to real-world DNN applications.
Among the defense strategies against AEs, adversarial training (AT) (Goodfellow et al., 2015; Madry et al., 2018) has been the state-of-the-art approach, which augments the training data with AEs to enhance robustness. However, it is widely recognized that AT-based methods suffer from a trade-off between robustness and accuracy (Tsipras et al., 2019): to achieve high robustness, they sacrifice accuracy on clean images. Despite the extensive number of studies on AT-based methods, this robustness-accuracy trade-off has been a huge obstacle to their practical use.
To this end, we investigate the following research question: “How can a model learn adversarially invariant representation without losing discriminative power?.” Specifically, we focus on representation-based invariance regularization for AT-based methods, in contrast to most of the existing defenses (Zhang et al., 2019; Wang et al., 2019) that employ invariance regularization on the predicted outputs.
@@
Our investigation reveals two key issues with naively applying invariance regularization (Fig.1(a)): (1) a “gradient conflict” between invariance loss and classification objectives, indicating the existence of “collapsing solutions,” and (2) the mixture distribution problem within Batch Normalization (BN) layers when the same BNs are used for both clean and adversarial inputs. “Gradient conflict” occurs when gradients for multiple loss functions oppose each other during joint optimization. Therefore, the observed conflict suggests that minimizing invariance loss may push the model toward non-discriminative directions. We hypothesize that it arises from the existence of “collapsing solutions” (Chen & He, 2021), i.e., a trivial solution for achieving invariance by encoding constant outputs. Furthermore, we find that using the same BNs for both clean and adversarial inputs causes the mixture distribution problem: the BNs are updated to have a mixture distribution of clean and adversarial inputs, which can be suboptimal for both clean and adversarial inputs.
To address these issues, we propose a novel method called Asymmetrically Representation-regularized Adversarial Training (AR-AT), incorporating a stop-gradient operation and a predictor, inspired by a modern non-contrastive self-supervised learning approach (Chen & He, 2021), and a split-BN structure, as depicted in Fig. 1(b). By employing the stop-gradient operation and predictor, AR-AT effectively avoids “collapsing solutions” and mitigates the “gradient conflict” between invariance loss and classification objectives during training. Additionally, the split-BN structure plays a key role in achieving high robustness and accuracy by resolving the mixture distribution problem. With both components combined, our method learns adversarially invariant representations without sacrificing discriminative power, leading to high robustness and accuracy.
Furthermore, we discuss the relevance of our findings to existing knowledge distillation (KD)-based defenses (Cui et al., 2021; Suzuki et al., 2023), which their relative successes had not been well understood. Specifically, we attribute their relative success to avoiding “collapsing solutions” by allowing variance between the teacher and student models and resolving the mixture distribution problem by using separate networks for clean and adversarial inputs.
@@
  • We rethink the representation-based invariance regularization in adversarial training to achieve high robustness without sacrificing accuracy.
  • We reveal two key issues hindering invariance regularization: (1) a “gradient conflict” between invariance loss and classification objectives, and (2) the mixture distribution problem within Batch Norm (BN) layers.
  • We propose a novel method AR-AT to incorporate a stop-gradient operation and a predictor, inspired by a recent non-contrastive self-supervised learning approach, and a split-BN structure.
  • Our method significantly improves the robustness-accuracy trade-off by effectively learning discriminative yet adversarially invariant representations.
  • We provide a new perspective on the relative success of knowledge distillation-based defenses, attributing it to avoiding “collapsing solutions” and resolving the mixture distribution problem.
@@ 2.1 Adversarial Attack
$x^{\prime}=\mathop{\rm arg~{}max}\limits_{\delta\in\mathcal{S}}\mathcal{L}(f_{\theta}(x+\delta),y)$ (1)
@@ 2.2 Adversarial Training
Adversarial training (AT) (Goodfellow et al., 2015; Madry et al., 2018), which augments the training data with adversarial examples (AEs), has been the state-of-the-art approach to defend against adversarial attacks. Originally, Goodfellow et al. (2015) proposed to train the model with AEs generated by the fast gradient sign method (FGSM); in contrast, Madry et al. (2018) proposed to train a model with much stronger AEs generated by PGD, which is an iterative version of FGSM. Formally, the standard AT (Madry et al., 2018) solves the following optimization problem:

### v3

@@ Abstract
Adversarial training often suffers from a robustness-accuracy trade-off, where achieving high robustness comes at the cost of accuracy. One approach to mitigate this trade-off is leveraging invariance regularization, which encourages model invariance under adversarial perturbations; however, it still leads to accuracy loss. In this work, we closely analyze the challenges of using invariance regularization in adversarial training and understand how to address them. Our analysis identifies two key issues: (1) a “gradient conflict" between invariance and classification objectives, leading to suboptimal convergence, and (2) the mixture distribution problem arising from diverged distributions between clean and adversarial inputs. To address these issues, we propose Asymmetric Representation-regularized Adversarial Training (AR-AT), which incorporates asymmetric invariance loss with stop-gradient operation and a predictor to avoid gradient conflict, and a split-BatchNorm (BN) structure to resolve the mixture distribution problem. Our detailed analysis demonstrates that each component effectively addresses the identified issues, offering novel insights into adversarial defense. AR-AT shows superiority over existing methods across various settings. Finally, we discuss the implications of our findings to knowledge distillation-based defenses, providing a new perspective on their relative successes.
@@ 1 Introduction
Computer vision models based on deep neural networks (DNNs) are vulnerable to adversarial examples (AEs) (Szegedy et al., 2014; Goodfellow et al., 2015), which are carefully perturbed inputs to fool DNNs. Since AEs can fool DNNs without affecting human perception, they pose potential threat to real-world DNN applications. Adversarial training (AT) (Goodfellow et al., 2015; Madry et al., 2018) has been the state-of-the-art defense strategy, however, AT-based methods suffer from a significant robustness-accuracy trade-off (Tsipras et al., 2019): to achieve high robustness, they sacrifice accuracy on clean images. Despite the extensive number of studies, this trade-off has been a huge obstacle to their practical use.
One of the promising approaches to mitigate this trade-off is to leverage invariance regularization, such as TRADES (Zhang et al., 2019) and LBGAT (Cui et al., 2021), which encourage the model to be invariant under adversarial perturbations. However, these methods still face trade-offs, and their limitations require further understanding.
To this end, we investigate the following research question: “How can a model learn adversarially invariant representations without compromising discriminative ability?." This work carefully analyzes the challenges of applying invariance regularization in AT to improve the robustness-accuracy trade-off. We identify novel issues and propose novel solutions to address them, offering novel insights into adversarial defense.
@@
Our investigation identifies two key issues in applying invariance regularization (Fig.1(a)): (1) a “gradient conflict" between invariance loss and classification objectives, leading to suboptimal convergence, and (2) the mixture distribution problem within Batch Normalization (BN) layers when the same BNs are used for both clean and adversarial inputs. “Gradient conflict" suggests that minimizing invariance loss may push the model toward non-discriminative directions and conflict with the classification objectives. Furthermore, we find that using the same BNs for both clean and adversarial inputs causes the mixture distribution problem: the BNs are updated to have a mixture distribution of clean and adversarial inputs, which can be suboptimal for both clean and adversarial inputs.
To address these issues, we propose a novel method, Asymmetric Representation-regularized Adversarial Training (AR-AT), incorporating an asymmetric invariance regularization with a stop-gradient operation and a predictor, and a split-BN structure, as depicted in Fig. 1(f). Our step-by-step analysis demonstrates that each component effectively addresses the identified issues, offering novel insights: stop-gradient operation resolves “gradient conflict," and split-BN structure resolves the mixture distribution problem. With both components combined, AR-AT improves the robustness-accuracy trade-off, outperforming existing methods across various settings.
Furthermore, we discuss the relevance of our findings to existing knowledge distillation (KD)-based defenses (Cui et al., 2021; Suzuki et al., 2023), pointing out the underlying factors of their effectiveness that have not yet been investigated. We attribute their effectiveness to resolving “gradient conflict” and the mixture distribution problem.
@@
  • We focus on understanding the challenges of invariance regularization in adversarial training to improve the robustness-accuracy trade-off. We identify novel issues and propose novel solutions.
  • We reveal two key issues in using invariance regularization: (1) a “gradient conflict" between invariance loss and classification objectives, and (2) the mixture distribution problem within Batch Norm (BN) layers.
  • We propose a novel method, AR-AT, which incorporates an asymmetric invariance regularization and a split-BN structure. Our detailed analysis shows that each component effectively adresses the identified issues, offering novel insights into adversarial defense.
  • AR-AT outperforms existing methods across various settings.
  • We present a new perspective on KD-based defenses, which have not been well understood.
@@ 2.1 Adversarial attack
$x^{\prime}=x+\delta,\;\;\text{where}\;\;\delta=\operatorname*{arg\,max}_{\delta\in\mathcal{S}}\mathcal{L}(f_{\theta}(x+\delta),y)$ (1)
@@ 2.2 Adversarial training
Adversarial training (AT), which augments the training data with AEs, has been the state-of-the-art approach to defend against adversarial attacks. Originally, Goodfellow et al. (2015) proposed to train the model with AEs generated by the fast gradient sign method (FGSM); in contrast, Madry et al. (2018) proposed to train a model with much stronger AEs generated by PGD, which is an iterative version of FGSM. Formally, the standard AT (Madry et al., 2018) solves the following optimization: